OpenStudy (anonymous):

Find an equation of the line passing through the two points. Write an equation in standard form. (2,0) and (-3,8)

OpenStudy (anonymous):

The slope or m = (y1-y2)/(x1-x2). That is, rise over run. So, m = (8-0)/(-3-2) = 8/(-5) = -8/5 so, the line so far is y= mx+b = -8/5x + b We plug in the first point: 0=-8/5 (2) + b solving for b: 8/5 * 2 = b (moved over to other side) 16/5 = b, so our final equation is: y = -8/5 x + 16/5 Checking: (plug in values from other point) 8 = -8/5 (-3) + 16/5 8 = 24/5 +16/5 = 40/5 = 8. Done!